hand sanitizer recipe has an H2O2 concentration of 0.88molL-1. What volume of 9.8molL-1 H2O2 would be required to make up 500mL of the hand sanitizer.


1
Expert's answer
2021-06-17T06:50:02-0400

molarity in sanitiser = 0.88 mol/l

molarity needed = 9.8 mol/l

volume of sanitiser = 500 ml


Moles of sanitiser needed in 1 ltr of sanitiser = 0.88


So, volume of H2O2 needed to make 1 ltr of solution = 0.88 * 22.4 ltr



Now we know that


M1V1 = M2V2


0.88 * 0.88 * 22.4 = 9.8 * V2


V2 = 1.77 ltr